OpenStudy (anonymous):

HELP PLEASE:( Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). y = 3x - 5 y = 3x - 2 y = 3x + 1 y = 3x + 3

OpenStudy (kira_yamato):

dy/dx = 3 y = 3x + C When x=1, y=-2 -2 = C+3 C = -5 Therefore the eqn is y = 3x - 5